# Does learning the right latent variables necessarily improve in-context learning?

### v1

@@ 3 Experiments
Surprisingly, and counter to our predictions above, we found that the explicit model provided no significant benefit in OOD settings. In synthetic regression tasks shown in Fig. 2 (a), all models failed to generalize and obtained substantially worse performance than during ID evaluation, with the implicit model actually being the one that had a slightly lower degradation in performance. In classification and compositional tasks shown in Fig. 2 (b, c), all models generalized fairly well OOD and with similar performance. In summary, explicit models appear to provide no benefit across our tasks, both in terms of ID and OOD performance.
@@
Figure 3: Performance comparisons on a subset of tasks where the true latent variable ${\bm{z}}$ and prediction function $g$ are known. Implicit models are in shown gray and explicit models with Transformer prediction are in blue. Models in purple are trained with an auxiliary loss to predict the true latent variable. Models in green are defined using the true prediction function. Using the known prediction function leads to significantly better OOD performance.
@@
Explicit models learn to infer the correct latent variable, but not how to use it. Why didn’t the explicit model provide any benefit? Our initial hypothesis was that the implicit model could be susceptible to learning shortcuts that are sufficient to reduce the training loss and easy to express using self-attention between the query ${\bm{x}}_{*}$ and context ${\mathcal{D}}$. By summarizing the context in a bottleneck ${\bm{z}}_{\psi}$, the explicit model should instead be forced to extract the true latent variable in order to minimize the training loss, thus learning a solution that is closer to the actual data-generating process. There are then two possible explanations for our results: (1) the explicit models did not learn to extract the true latent variable despite inductive biases to do so induced by the bottleneck, or (2) they did extract the true latent variable but did not learn to use it in the correct way. We performed several experiments to test these two possibilities, and found strong evidence for the second.
@@
To test whether or not the explicit models failed because they did not extract the correct latent variable, we first attempted to encourage them to do so more directly by training an additional linear model which took ${\bm{z}}_{\psi}$ as input and attempted to predict the true latents ${\bm{z}}$. The loss of this linear model was then backpropagated through the context model along with the task loss. Results in Fig. 3 (purple) show that this auxiliary loss provided no improvement apart from minor increases in accuracy on Raven’s PM, suggesting that incorrect latent variable inference does not explain the explicit model’s suboptimal performance. Indeed, when we did not use the auxiliary loss as a training signal for the explicit model and just trained separate linear readouts on ${\bm{z}}_{\psi}$ using a stop-gradient, we found that we could still accurately decode the true latent variable from it (see Fig. 4 (a)). This means that in the absence of any explicit training signal to predict the true latent variable, the explicit model still learns to extract it in order to minimize prediction loss on the query.
Figure 4: Explicit models are interpretable as the bottleneck allows us to (a) linearly decode the true latent, and (b) intervene on it to obtain correct counterfactual predictions. Implicit models are shown in gray, explicit models with Transformer prediction in blue, and with MLP prediction in orange. Models in green use the true prediction function $g$. To evaluate decoding performance in (a), we also build a baseline by linearly decoding the true latent directly from concatenated context examples, which yielded significantly worse performance than decoding from the bottleneck. Baseline performances in units of the plots are – Linear regression: $0.49$, Nonlinear regression (MLP): $0.94$, Sinusoid regression: $0.33$, Linear classification: $0.86$, Nonlinear classification (MLP): $0.97$, Raven’s PM: $0.5$, and Gene targeting: $0.0$. In (b), the “Relative accuracy” takes in account the baseline (details in C.4).
@@
Given that the explicit model correctly infers the true latent variable in its bottleneck, we study whether the prediction function is suboptimally learned. In other words, despite the explicit model having access to the true ${\bm{z}}$, we hypothesize that $p_{\gamma}(y_{*}|{\bm{x}}_{*},{\bm{z}}_{\psi})$ does not learn the true data-generating process $y_{*}=g({\bm{x}}_{*},{\bm{z}})$, where $g$ is the true prediction function – e.g., for linear regression $g({\bm{x}}_{*},{\bm{z}})={\bm{z}}^{T}{\bm{x}}_{*}$. To validate this hypothesis, we trained explicit models with the prediction function $g$ hard-coded as an oracle. For instance, in the linear regression task, the ${\bm{z}}_{\psi}$ output by the context model was linearly projected to the same dimensionality as the true weights ${\bm{z}}$, after which the prediction function took its dot product with queried input ${\bm{x}}_{*}$. In this setting, if the explicit model extracts the correct latent variable, it should generalize perfectly both in and out of distribution. Our results in Fig. 3 confirm that using the correct prediction function indeed provides substantially better OOD generalization and effectively solves most tasks. This finding has significant implications: it suggests that while learning the true task latents may be a necessary condition for generalization, this must also be supplemented with significant inductive biases in the prediction function – for instance, through novel architectures – so that these task latents can be leveraged correctly. In the absence of such inductive biases, inferring the correct task latent appears to provide no benefits in practice. We do note that our nonlinear regression tasks, where ${\bm{z}}$ represents the weights of an underlying MLP generating the data, were an exception to the results described here in that using an oracle prediction function performed poorly. In this case, we conjecture that the underlying latent variable is too difficult to accurately infer from the context, while shortcut-based solutions would avoid latent variable inference altogether to provide robust solutions.
@@
Explicit models are easily interpretable. While explicit models do not provide performance benefits, the ability to extract the correct latent and summarize it in a single bottleneck can still be useful for interpretability. On tasks with a known underlying latent variable, we were able to linearly decode it from ${\bm{z}}_{\psi}$ with high accuracy in most cases, meaning that the information is not only present but also easily accessible (Fig. 4 (a)). The exceptions were on complex nonlinear regression tasks where the latents are difficult to infer and classification tasks where more context samples are needed to precisely identify the true decision boundaries. In contrast, finding such clear task-relevant latents is immensely challenging in an ordinary Transformer trained to do ICL, given that they can be distributed across a mixture of many layers and token positions.
@@
Furthermore, even when latent variables appear to be successfully identified using a linear decoder in some hidden layer of a Transformer, one often finds that the relationships merely amount to spurious correlations (Ravichander et al., 2021). For instance, if one manually modifies the activations in this hidden layer such that a different latent variable is predicted by the linear decoder, the model’s predictions do not generally change in a way that is “counterfactually” consistent with this new latent (i.e., the prediction is not what it should have been under the new latent variable). We therefore used the Distributed Alignment Search (DAS) method from Geiger et al. 2023b (see §C.4) to search for units in the implicit and explicit models that can be manipulated to obtain correct counterfactual predictions. For the explicit model, we limited this search to the bottleneck ${\bm{z}}_{\psi}$. We found that using the explicit model, we were indeed able to manipulate ${\bm{z}}_{\psi}$ and obtain correct counterfactual predictions, but we were not able to successfully do this using the implicit model, as shown in Fig. 4 (b). Explicit models might therefore be useful for both mechanistic interpretability and scientific discovery (Geiger et al., 2023a), where we do not know the underlying task latents and want to be able to easily uncover them from the trained model, and subsequently obtain a good predictor for an intervened system zero-shot given some knowledge about the intervention.
Figure 5: We analyze (a) Linear regression OOD task performance and (b) latent variable linear decoding performance as a function of model and task parameters. Task performance scales similarly for implicit (gray) and explicit models with Transformer prediction (blue). Models in green use the true prediction function $g$. Latent variable decoding accuracy scales as a function of the latent variable’s inherit uncertainty (data dimensionality and context length) and the model’s size.
@@
Scaling Trends across Different Properties. To better compare the implicit and explicit models, we investigated their OOD task performance on linear regression as we varied the different properties of the task (input dimensionality and context length) and the size of the model used (Fig. 5 (a)). We found that performance in both models scaled similarly, but that the implicit model reliably outperformed the explicit one unless it used the known prediction function $g$. We also looked at the latent variable linear decoding accuracy in the explicit model as a function of these task and model properties (Fig. 5 (b)). As expected, we found that the latent variable was easier to decode from the explicit model’s bottleneck when there was less inherit uncertainty about its value (lower data dimensionality, longer context length) and when the explicit model was given more capacity. However, throughout the different settings, we see that while the explicit model does learn the true latent well, it is not sufficient to get a performance boost over the implicit models. Further details on the setup of these scaling experiments is provided in §C.2.
4 Related Work
@@
In-Context Learning. In-Context learning (ICL) is an ability of certain trained models to take an entire task’s dataset as input and parameterize solutions directly in their layer activations, which then condition subsequent computation on novel inputs from those same tasks. Generally, this ability is found in sequence models such as Transformers where the task dataset, or “context”, corresponds to an earlier part of the sequence. ICL was first observed in large-scale pre-trained LLMs (Brown et al., 2020), and is similar in many respects to meta-learning (Chan et al., 2022) (see §A). These LLM findings were subsequently expanded to more controlled settings outside of the language modality, where Transformer models were directly trained on task distributions such as linear regression (Von Oswald et al., 2023; Garg et al., 2023), hidden Markov models (Xie et al., 2022), compositional grammars (Hahn & Goyal, 2023), regular languages (Akyürek et al., 2024) and Turing machines (Grau-Moya et al., 2024), with a set of task observations defining the “context”. These works highlight that Transformers are indeed able to model many types of complex task distributions, approaching in many cases the performance of the Bayes-optimal predictor, the a-priori optimal solution (Xie et al., 2022). Our work lies along similar lines but using more complex tasks and a systematic study into the differences between modeling the predictive space directly, or through a two-step process involving explicit inference of task latents.
@@
Shortcuts in ICL. Shortcut learning is a phenomenon that has widely been observed in machine learning (Geirhos et al., 2020), and refers to where a model solves a task through statistical correlations that are accidental and thus not robust to even slight distribution shifts. A classical example of this in image classification is the usage of background cues to classify objects (Ribeiro et al., 2016). Similar mistakes are know to be very common in NLP (McCoy et al., 2019), specifically in reasoning tasks (Zhang et al., 2022). Particularly relevant to our work, many authors have shown that has shown that Transformers are very prone to relying on shortcuts when performing ICL Tang et al. (2023). For instance, Olsson et al. (2022); Singh et al. (2024) have shown that induction heads play in important role in ICL by predicting that the continuation of a token will be the same as last time (i.e. $[a][b]\ldots[a]\rightarrow[b]$). As shown by Von Oswald et al. (2023) this motif can be used to do linear regression, and can generally be seen as a form of kernel regression (i.e. $p(y_{q}|x_{q},x_{1:n})\propto\sum_{i}K(x_{i},x_{q})y_{i}$, (Han et al., 2023)). This observation draws a link between those types of solutions and non-parametric inference methods in statistics (Wikipedia, 2024), of which kernel regression is a member. In contrast (Hendel et al., 2023; Todd et al., 2023) have concurrently shown that in some cases Transformers encode a “task vector” that they infer from the context and then use to do the prediction. There is therefore a need to better understand the nature of shortcuts in ICL (which relates to the field of mechanistic interpretability, see §A) and whether or not they can be easily avoided for better generalization. Our work explores this very question.
@@
5 Conclusion
A commonly believed hypothesis is that Transformers do ICL through brittle statistical shortcuts rather than by inferring the underlying generative latent variables of the task, and that this explains their inability to generalize outside of the training distribution. Here, we empirically tested this hypothesis by minimally modifying the Transformer architecture through the use of a bottleneck that factorized the model into separate context aggregation and prediction functions, creating an inductive bias for explicit latent variable inference. While we confirmed that this model indeed learned to infer the correct latent variables across many ICL tasks, it surprisingly gave no improvement in performance for either in-distribution or out-of-distribution evaluation. Contrary to common belief, then, we showed that simply learning the correct latent variables for the tasks is not sufficient for better generalization because end-to-end optimization does not learn the right prediction model to leverage these latent variables. Indeed, when we substituted the prediction function with a known oracle, we often saw significant advantages in OOD performance. Importantly, we were able to make such findings because we performed controlled experiments on simple tasks with known latent variables and prediction functions, thus avoiding the complexities and confounding variables that come with natural language and pre-trained LLMs. Our results point to a line of future work that incorporates inductive biases in the prediction model to better leverage the inferred latent variables, and motivates current efforts in improving (a) amortized methods for in-context prediction (e.g., Garnelo et al., 2018a, b; Kim et al., 2019; Hollmann et al., 2022) and (b) neurosymbolic AI which tries to combine flexible LLMs for inference with highly structured functions for prediction (e.g., Wong et al., 2023; Mialon et al., 2023; Ellis, 2023).
Limitations. While our work investigates potential reasons behind the shortcomings of current ICL approaches, overcoming those shortcomings is an entirely different and bigger problem which is a relevant future direction but not an achievement of the current work. We also look at ICL solely from the perspective of algorithmic tasks with models trained from scratch, without exploring in detail the connections with ICL in pre-trained models.
Broader Impact. Our approach paves the way into an interpretable model for in-context learning, which could prevent harmful use of such models by additionally encoding constraints on the interpretable bottleneck. Additionally, our work provides a better understanding about ICL which can lead to safer future models.
Acknowledgments and Disclosure of Funding
The authors would like to thank Tom Marty, Johannes von Oswald, Joao Sacramento and Yoshua Bengio for discussions. The authors would like to acknowledge the computing resources provided by the Mila cluster to enable the experiments outlined in this work. SM acknowledges the support of UNIQUE’s scholarship. EE acknowledges the support of the Vanier CGS scholarship. GL and DS acknowledge the support of the Canada CIFAR AI Chair program and Google Research. GL is also supported by NSERC Discovery Grant RGPIN-2018-04821, and a Canada Research Chair in Neural Computations and Interfacing. The authors also thank NVIDIA for computing resources.
Appendix A Additional Related Work
@@
Meta-Learning. Meta-learning (Weng, 2018; Hospedales et al., 2020) studies systems that can learn over two levels: rapidly through an inner-loop that is meta-learned using a slower outer-loop. The goal in such methods is to learn a good initialization common to the parameterized family of tasks, in a manner that obtaining a particular solution for a new task is fast from this initial point. The inner loop provides an optimization trajectory for a randomly drawn task from some initialization, which in itself is optimized in the outer loop to a good solution applicable for the global set of tasks. Typically, evaluation is done on some meta-validation set of tasks not seen during training. Task distributions can for example be a set of different classification/regression tasks (few shot learning, Vettoruzzo et al. (2023)) or variations of a reinforcement learning (meta-RL, Beck et al. (2023)). The goal is similar to ICL approaches in the sense that given a novel context ${\mathcal{D}}$, one wants to make predictions for some query ${\bm{x}}_{*}$. However, a big difference is that ICL approaches bypass modeling a common initialization by working directly on the prediction side (implicit), or instead predict the optimal parameters directly zero-shot through inference on the context model (explicit).
@@
Mechanistic Interpretability. Mechanistic interpretability is interested in understanding deep neural network’s computations through interpretable abstraction, akin to what computational neuroscience does with the brain. Alain & Bengio (2018) introduced the foundational technique of linear “probes”, which are linear models trained on the hidden state of a network to predict a abstract feature of the input; the success of which would suggest that such feature is used by the model. Since then, this naïve approach has been criticized for being potentially misleading (Ravichander et al., 2021); in many cases a feature can be linearly decoded from a model without the model using it. More reliable methods grounded in causality (Vig et al., 2020; Geiger et al., 2024) have now became the gold standard, and their use applied to Transformers have been exploding in popularity in the recent years (Elhage et al., 2021). In this paper we use linear probing for its simplicity for most experiments but use more sophisticated on one experiment to show that the learned feature are causally meaningful.

### v2

@@ 4 Experiments
Surprisingly, and counter to our predictions above, we found that the explicit model provided no significant benefit in OOD settings. In synthetic regression tasks shown in Figure 2 (a), all models failed to generalize and obtained substantially worse performance than during ID evaluation, with the implicit model actually being the one that had a slightly lower degradation in performance. In classification and compositional tasks shown in Figure 2 (b, c), all models generalized fairly well OOD and with similar performance. Our results on reusable modular MoE task in Figure 5 further indicate that implicit models consistently outperform explicit ones across different proportion of latent combinations seen during training (X-axis). In summary, explicit models appear to provide no benefit across our tasks, both in terms of ID and OOD performance.
@@
Explicit models learn to infer the correct latent variable, but not how to use it. Why didn’t the explicit model provide any benefit? Our initial hypothesis was that the implicit model could be susceptible to learning shortcuts that are sufficient to reduce the training loss and easy to express using self-attention between the query ${\bm{x}}_{*}$ and context ${\mathcal{D}}$. By summarizing the context in a bottleneck ${\bm{z}}_{\psi}$, the explicit model should instead be forced to extract the true latent variable in order to minimize the training loss, thus learning a solution that is closer to the actual data-generating process. There are then two possible explanations for the results in Figure 2: (1) the explicit models did not learn to extract the true latent variable despite inductive biases to do so induced by the bottleneck, or (2) they did extract the true latent variable but did not learn to use it in the correct way. We performed several experiments to test these two possibilities, and found strong evidence for the second.
@@
To test whether or not the explicit models failed because they did not extract the correct latent variable, we added an additional supervised loss term to Equation 1, $||{\bm{z}}-W{\bm{z}}_{\psi}||^{2}$, to force the explicit model to recover the true task latent ${\bm{z}}$ up to a linear transformation, where $W$ is a learnable parameter. The loss of this linear model was then backpropagated through the context model along with the task loss. Results in Figure 3 (purple) show that this auxiliary loss provided no improvement apart from minor increases in accuracy on Raven’s PM, suggesting that incorrect latent variable inference does not explain the explicit model’s suboptimal performance. Indeed, when we did not use the auxiliary loss as a training signal for the explicit model and just evaluated the quality of the learned task latents by training a separate linear predictor to predict ${\bm{z}}_{\psi}$, we found that we could still accurately predict the true latent variable (see Figure 4 (a)). This means that in the absence of any explicit training signal to predict the true task latent, the explicit model accurately learns task latents up to linear reparameterizations.
Figure 7: We analyze sinusoid regression by sampling the query either from the normal distribution (left) or close to the context (right) during training but always far from context during evaluation. We see that when queries are sampled close to the context, implicit models which can rely more on kernel-regression based nearest-neighbor solutions don’t generalize far from context, while explicit models do.
@@
Given that the explicit model correctly infers the true latent variable in its bottleneck, we study whether the prediction function is suboptimally learned. In other words, despite the explicit model having access to the true ${\bm{z}}$, we hypothesize that $p_{\gamma}(y_{*}|{\bm{x}}_{*},{\bm{z}}_{\psi})$ does not learn the true data-generating process $y_{*}=g({\bm{x}}_{*},{\bm{z}})$, where $g$ is the true prediction function – e.g., for linear regression $g({\bm{x}}_{*},{\bm{z}})={\bm{z}}^{T}{\bm{x}}_{*}$. To validate this hypothesis, we trained explicit models with the prediction function $g$ hard-coded as an oracle. For instance, in the linear regression task, the ${\bm{z}}_{\psi}$ output by the context model was linearly projected to the same dimensionality as the true weights ${\bm{z}}$, after which the prediction function took its dot product with queried input ${\bm{x}}_{*}$. In this setting, if the explicit model extracts the correct latent variable, it should generalize perfectly both in and out of distribution. Our results in Figure 3 confirm that using the correct prediction function indeed provides substantially better OOD generalization and effectively solves most tasks. This finding has significant implications: it suggests that while learning the true task latents may be a necessary condition for generalization, this must also be supplemented with significant inductive biases in the prediction function – for instance, through novel architectures – so that these task latents can be leveraged correctly. In the absence of such inductive biases, inferring the correct task latent appears to provide no benefits in practice. We do note that our nonlinear regression tasks, where ${\bm{z}}$ represents the weights of an underlying MLP generating the data, were an exception to the results described here in that using an oracle prediction function performed poorly. In this case, we conjecture that the underlying latent variable is too difficult to accurately infer from the context, while shortcut-based solutions would avoid latent variable inference altogether to provide robust solutions.
@@
Explicit models are easily interpretable. While explicit models do not provide performance benefits, the ability to extract the correct latent and summarize it in a single bottleneck can still be useful for interpretability. On tasks with known underlying latent variables, we were able to linearly decode them from ${\bm{z}}_{\psi}$ with high accuracy, meaning that the information is not only present but also easily accessible (Figure 4 (a)). The exceptions were complex nonlinear regression tasks where the latents are difficult to infer and classification tasks where more context samples are needed to precisely identify the decision boundary. In contrast, finding such clear task-relevant latents is immensely challenging in an ordinary Transformer trained to do ICL, given that they can be distributed across a mixture of many layers and token positions.
@@
Furthermore, even when latent variables appear to be successfully identified using a linear decoder in some hidden layer of a Transformer, one often finds that the relationships merely amount to spurious correlations (Ravichander et al., 2021). For instance, if one manually modifies the activations in this hidden layer such that a different latent variable is predicted by the linear decoder, the model’s predictions do not generally change in a way that is “counterfactually” consistent with this new latent (i.e., the prediction is not what it should have been under the new latent variable). We therefore used the Distributed Alignment Search (DAS) method from Geiger et al. 2023b (see subsection C.4) to search for units in the implicit and explicit models that can be manipulated to obtain correct counterfactual predictions. For the explicit model, we limited this search to the bottleneck ${\bm{z}}_{\psi}$. We found that using the explicit model, we were indeed able to manipulate ${\bm{z}}_{\psi}$ and obtain correct counterfactual predictions, but we were not able to successfully do this using the implicit model, as shown in Figure 4 (b). Explicit models might therefore be useful for both mechanistic interpretability and scientific discovery (Geiger et al., 2023a), where we do not know the underlying task latents and want to be able to easily uncover them from the trained model, and subsequently obtain a good predictor for an intervened system zero-shot given some knowledge about the intervention.
@@
Scaling Trends across Different Properties. To better compare the implicit and explicit models, we investigated their OOD task performance on linear regression as we varied the different properties of the task (input dimensionality and context length) and the size of the model used (Figure 6 (a)). We found that performance in both models scaled similarly, but that the implicit model reliably outperformed the explicit one unless it used the known prediction function $g$. We also looked at the latent variable linear decoding accuracy in the explicit model as a function of these task and model properties (Figure 6 (b)). As expected, we found that the latent variable was easier to decode from the explicit model’s bottleneck when there was less inherit uncertainty about its value (lower data dimensionality, longer context length) and when the explicit model was given more capacity. However, throughout the different settings, we see that while the explicit model does learn the true latent well, it is not sufficient to get a performance boost over the implicit models. Further details on the setup of these scaling experiments is provided in subsection C.2.
Figure 8: We compare using the auxiliary ground-truth task latent loss directly on the output of context aggregation, i.e. ${\bm{z}}_{\psi}({\mathcal{D}})$ (aux_direct), or to a linear decoding from it (aux_decoding).
Impact of auxiliary loss on decoding from bottleneck. Additionally, we perform an experiment where instead of using an auxiliary loss obtained between the ground-truth task latents ${\bm{z}}$ and a decoding from the explicit model’s bottleneck $||{\bm{z}}-W{\bm{z}}_{\psi}||^{2}$ (called aux_decoded), we instead force the bottleneck itself to be directly close to the ground-truth $||{\bm{z}}-{\bm{z}}_{\psi}||^{2}$ (called aux_direct). Since the prediction function relies on the bottleneck and not its decoding, removing this extra layer when providing additional supervision might allow the bottleneck to better reflect the task latents and thereby aid prediction. Our results, however, indicate that doing so does not lead to any benefits on OOD evaluation for linear regression, further strengthening the conclusion that effective task latent inference is not the biggest problem in such models.
Extreme Shortcut Injection. Finally, we test whether injecting extreme shortcuts during training pushes implicit models to learn nearest-neighbor styled kernel-regression solutions as opposed to uncovering the underlying functional form. We consider two cases, where queries during training are sampled (a) randomly, or (b) near context points. The latter further incentivizes implicit models to learn nearest neighbour shortcuts. At evaluation, the queries are sampled far from the context. Our results on sinusoid regression in Figure 7 indicate that while implicit models perform well generally, they suffer considerably more in the presence of such injected shortcuts since explicit models distill the task latent from context independent of the query.
We further refer to Appendix D for a detailed analysis.
5 Conclusion
A commonly believed hypothesis is that Transformers do ICL through brittle statistical shortcuts rather than by inferring the underlying generative latent variables of the task, and that this explains their inability to generalize outside of the training distribution. Here, we empirically tested this hypothesis by minimally modifying the Transformer architecture through the use of a bottleneck that factorized the model into separate context aggregation and prediction functions, creating an inductive bias for explicit latent variable inference. While we confirmed that this model indeed learned to infer the correct latent variables across many ICL tasks, it surprisingly gave no improvement in performance for either in-distribution or out-of-distribution evaluation. Contrary to common belief, then, we showed that simply learning the correct latent variables for the tasks is not sufficient for better generalization because end-to-end optimization does not learn the right prediction model to leverage these latent variables.
Impact Statement
This paper provides a comparative analysis to better understand the capabilities of current in-context learners from a point of view of making them more robust and aligned with the true underlying models of the data. We believe that this is an important step towards understanding when and how machine learning models can rely on shortcuts and spurious correlations, and understanding whether such correlations can be mitigated through the use of conditional independence assumptions and bottlenecks, as is investigated in this work.
While we show a negative result that such bottlenecks do not substantially aid generalization, they do come with interpretability benefits which are extremely useful when deploying AI systems at scale. Finally, we hope that our analysis sparks controlled experiments to understand the mechanisms behind in-context learning better as well as incorporating and validating numerous inductive biases to see whether they do aid generalization and reduce the reliance on shortcuts.
Ackonwledgements
The authors would like to acknowledge the following researchers for valuable discussions and exchanges: Joao Sacramento, Johannes von Oswald. All authors acknowledge support from an unrestricted gift from Google inc. EE acknowledges support from Vanier Canada Graduate Scholarship #492702. SM acknowledges the support of PhD Excellence Scholarship from UNIQUE. DS acknowledges support from NSERC Discovery Grant RGPIN-2023-04869, and a Canada-CIFAR AI Chair. GL acknowledges support from NSERC Discovery Grant RGPIN-2018-04821, the Canada Research Chair in Neural Computations and Interfacing, and a Canada-CIFAR AI Chair.
Appendix A Related Work
@@
In-Context Learning. In-Context learning (ICL) is an ability of certain trained models to take an entire task’s dataset as input and parameterize solutions directly in their layer activations, which then condition subsequent computation on novel inputs from those same tasks. Generally, this ability is found in sequence models such as Transformers where the task dataset, or “context”, corresponds to an earlier part of the sequence. ICL was first observed in large-scale pre-trained LLMs (Brown et al., 2020), and is similar in many respects to meta-learning (Chan et al., 2022). These LLM findings were subsequently expanded to more controlled settings outside of the language modality, where Transformer models were directly trained on task distributions such as linear regression (Von Oswald et al., 2023; Garg et al., 2023), hidden Markov models (Xie et al., 2022), compositional grammars (Hahn & Goyal, 2023), regular languages (Akyürek et al., 2024) and Turing machines (Grau-Moya et al., 2024), with a set of task observations defining the “context”. These works highlight that Transformers are indeed able to model many types of complex task distributions, approaching in many cases the performance of the Bayes-optimal predictor, the a-priori optimal solution (Xie et al., 2022). Our work lies along similar lines but using more complex tasks and a systematic study into the differences between modeling the predictive space directly, or through a two-step process involving explicit inference of task latents.
@@
Shortcuts in ICL. Shortcut learning is a phenomenon that has widely been observed in machine learning (Geirhos et al., 2020), and refers to where a model solves a task through statistical correlations that are accidental and thus not robust to even slight distribution shifts. A classical example of this in image classification is the usage of background cues to classify objects (Ribeiro et al., 2016). Similar mistakes are know to be very common in NLP (McCoy et al., 2019), specifically in reasoning tasks (Zhang et al., 2022). Particularly relevant to our work, many authors have shown that has shown that Transformers are very prone to relying on shortcuts when performing ICL (Tang et al., 2023a). For instance, (Olsson et al., 2022; Singh et al., 2024) have shown that induction heads play in important role in ICL by predicting that the continuation of a token will be the same as last time (i.e. $[a][b]\ldots[a]\rightarrow[b]$). As shown by (Von Oswald et al., 2023) this motif can be used to do linear regression, and can generally be seen as a form of kernel regression (i.e. $p(y_{q}|x_{q},x_{1:n})\propto\sum_{i}K(x_{i},x_{q})y_{i}$, (Han et al., 2023)). This observation draws a link between those types of solutions and non-parametric inference methods in statistics (Hastie, 2009), of which kernel regression is a member. In contrast (Hendel et al., 2023; Todd et al., 2023) have concurrently shown that in some cases Transformers encode a “task vector” that they infer from the context and then use to do the prediction. There is therefore a need to better understand the nature of shortcuts in ICL and whether or not they can be easily avoided for better generalization. Our work explores this very question.
@@
Meta-Learning. Meta-learning (Hospedales et al., 2020b, a) studies systems that can learn over two levels: rapidly through an inner-loop that is meta-learned using a slower outer-loop. The goal in such methods is to learn a good initialization common to the parameterized family of tasks, in a manner that obtaining a particular solution for a new task is fast from this initial point. The inner loop provides an optimization trajectory for a randomly drawn task from some initialization, which in itself is optimized in the outer loop to a good solution applicable for the global set of tasks. Typically, evaluation is done on some meta-validation set of tasks not seen during training. Task distributions can for example be a set of different classification/regression tasks (few shot learning, (Vettoruzzo et al., 2023)) or variations of a reinforcement learning (meta-RL, (Beck et al., 2023)). The goal is similar to ICL approaches in the sense that given a novel context ${\mathcal{D}}$, one wants to make predictions for some query ${\bm{x}}_{*}$. However, a big difference is that ICL approaches bypass modeling a common initialization by working directly on the prediction side (implicit), or instead predict the optimal parameters directly zero-shot through inference on the context model (explicit).
@@
Mechanistic Interpretability. Mechanistic interpretability is interested in understanding deep neural network’s computations through interpretable abstraction, akin to what computational neuroscience does with the brain. (Alain & Bengio, 2018) introduced the foundational technique of linear “probes”, which are linear models trained on the hidden state of a network to predict an abstract feature of the input; the success of which suggests that such a feature is used by the model. Since then, this naïve approach has been criticized for being potentially misleading (Ravichander et al., 2021); in many cases a feature can be linearly decoded from a model without the model using it. More reliable methods grounded in causality (Vig et al., 2020; Geiger et al., 2024) have now became the gold standard, and their use applied to Transformers has been exploding in popularity (Elhage et al., 2021).
Figure 9: Same experiment as Figure 4b but with the linear regression task. Specifically, we use Distributed Alignment Search (DAS, (Geiger et al., 2023b), see Appendix C for details) to find the 10 dimensional subspace in each model with best simulates counterfactual interventions on the task vector (in this case, the weight of the linear regression). In both explicit model, the subspace is taken at the bottleneck. In the implicit one, we perform the DAS at all layer of the query token and report the best one. The reported metric is the MSE of the intervened model on the intervened regression problem (i.e. using the same query $x$ but $y$’s coming from the intervened linear regression weights, Appendix C for details).